3. You have a couple of charged spheres. Sphere A has a charge of 3.4C and a mass of 2.3 grams. The second sphere (B) has a charge of 2.3C and a mass of 5.6 grams. They are initially placed 1.2cm apart. When they are released what is each of the spheres initial accelerations?
Answer #1

Sphere A: mass = 2.3 gm =0.0023 kg, charge = 3.4 C

Sphere B : mass = 5.6 gm = 0.0056 kg, charge = 2.3 C

distance between two charges = 1.2 cm = 0.012 m

force F= (k x QA x QB)/(d2) = (9 x 109 x 3.4 x 2.3) / (0.012)2 = 7.038 x 1010 / 1.44 x 10-4 = 4.8875 x 1014 N

F = Mass X acceleration

Acceleration = Force / mass

Acceleration of sphere A = 4.8875 x 1014 / 0.0023 = 2.125 x 1017 m/s2

Acceleration of sphere B = 4.8875 x 1014 / 0.0056 = 8.727 x 1016 m/s2